# DiMA: An LLM-Powered Ride-Hailing Assistant at DiDi

### v1

@@ Abstract.
On-demand ride-hailing services like DiDi, Uber, and Lyft have transformed urban transportation, offering unmatched convenience and flexibility. In this paper, we introduce DiMA, an LLM-powered ride-hailing assistant deployed in DiDi Chuxing. Its goal is to provide seamless ride-hailing services and beyond through a natural and efficient conversational interface under dynamic and complex spatiotemporal urban contexts. To achieve this, we propose a spatiotemporal-aware order planning module that leverages external tools for precise spatiotemporal reasoning and progressive order planning. Additionally, we develop a cost-effective dialogue system that integrates multi-type dialog repliers with cost-aware LLM configurations to handle diverse conversation goals and trade-off response quality and latency. Furthermore, we introduce a continual fine-tuning scheme that utilizes real-world interactions and simulated dialogues to align the assistant’s behavior with human prefered decision-making processes. Since its deployment in the DiDi application, DiMA has demonstrated exceptional performance, achieving $93\%$ accuracy in order planning and $92\%$ in response generation during real-world interactions. Offline experiments further validate DiMA’s capabilities, showing improvements of up to $70.23\%$ in order planning and $321.27\%$ in response generation compared to three state-of-the-art agent frameworks, while reducing latency by 0.72×0.72$\times$0.72 × to 5.47×5.47$\times$5.47 ×. These results establish DiMA as an effective, efficient, and intelligent mobile assistant for ride-hailing services. Refer to caption Figure 1. The conversational interface in DiDi Chuxing App. Given a user query issued on 2024-08-28, (a) DiMA proactively guides the user to complete necessary spatiotemporal trip information, and (b) automatically schedules a future ride-hailing order from Guangdong Museum East Gate to Terminal T2 of Guangzhou Baiyun International Airport.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generation
@@ 1. Introduction
On-demand ride-hailing services like DiDi, Uber, and Lyft have significantly transformed urban transportation systems and human mobility patterns (Xu et al., 2018). These platforms have introduced a paradigm shift in how people access transportation, offering unparalleled convenience, efficiency, and flexibility (Feng et al., 2021; Liu et al., 2022). As one of the largest ride-hailing platforms, the DiDi mobile application processes over ten billion user requests daily, encompassing tasks such as destination selection, travel time estimation, ride-hailing, re-routing, and feedback. These interactions span all stages of the ride-hailing process including pre-ride, in-ride, and post-ride (S. Alpers et al., 2020). Despite their transformative impact, existing mobile applications still rely heavily on manual inputs like typing, clicking, and scrolling, making the process time-intensive and often inefficient (Hsu and Chen, 2023).
@@
However, building an LLM-based ride-hailing assistant is a non-trivial task due to the following three challenges. (1) Spatiotemporal Travel Intention Understanding. The ride-hailing process relies heavily on understanding the spatiotemporal trip information such as destination and departure time. LLMs often struggle with spatial reasoning (Xie et al., 2024) and temporal calculations (Su et al., 2024). For instance, when issuing a query shown in Figure 1, general LLMs may fail to derive the exact date of next Friday and neglect the user’s current location as the start location. As reported in Figure 2(a), our quantitative studies validate that even state-of-the-art LLMs like GPT-4o and Qwen2.5-72B achieve less than 80% accuracy in spatiotemporal travel intention understanding, therefore limiting their applicability for handling real-world ride-hailing requests. (2) Proactive Order Planning. Ride-hailing involves multiple factors (e.g., waiting time, willing car type, destination) for matching passengers and drivers, which may require multi-turn conversations. Moreover, during a conversation, users may cancel confirmed orders, request orders with infeasible conditions (e.g., unrealistically low costs), or engage in ride-hailing policy inquiry. How to proactively and naturally guide users through the order planning process also poses a significant challenge. (3) Cost-Effective Response Generation. The assistant should generate responses adapted to varying conversation goals and order statuses. For instance, guiding users to place an order often requires formulaic language responses, handling infeasible request needs clarification, whereas answering ride-hailing policy queries may involve retrieving DiDi’s knowledge base. As shown in Figure 2(b), in general, larger LLMs tend to be more versatile to handle complex scenarios but result in higher response latency. Generating accurate responses within strict service latency is another critical challenge.
@@
Since May 2024, DiMA has been deployed in the DiDi application, as depicted in Figure 1. Two months’ online experiments in production demonstrate that DiMA achieves 93% accuracy in order planning and 92% in response generation. We further conduct offline experiments on synthetic datasets in Beijing and Shanghai. The results validate that DiMA significantly outperforms three state-of-the-art agent frameworks on order planning and response generation tasks, with improvements up to 70.23% and 321.27%, respectively. Moreover, the results show that DiMA maintains comparable performance to models like Qwen2.5-72B, Mistral-123B, and Llama3.1-405B, while reducing latency by 0.72x to 5.47x.
Our contributions are summarized as follows:
  • We propose the first LLM-based assistant deployed in DiDi mobile application, offering a new interaction alternative for ride-hailing.
  • We develop a spatiotemporal-aware tool-augmented order planning module and a cost-effective dialogue system to enhance accuracy and efficiency.
  • We introduce a continual fine-tuning scheme to ensure sustained proactive order planning guidance aligned with human decision processes.
  • Extensive experiments validate DiMA’s effectiveness and demonstrate its advantages against state-of-the-art LLM agent frameworks in both online and offline settings.
@@ 2. Preliminaries
In this section, we first present the formal problem definition, and then briefly introduce the evolution of DiDi’s ride-hailing service.
@@ 4.2. Temporal Reasoning
Similar to spatial mapping, we construct a four stage pipeline: (1) Temporal tool collection and annotation, (2) Temporal decomposition. (3) Tool selection, and (4) Tool calling. The whole process is similar with spatial mapping. Take the user query in Figure 1 again for example. Since the departure time is not explicitly specified, we follow product design to threat mentioned time as departure time. Therefore, DiMA automatically generate structured function call Get_departure_time() with key parameters {”date”: ”1 week later, Fri”, ”time”: ”18:00:00”} to obtain user’s departure time. Then, based on the issued date ”2024-08-28”, DiMA calls annotated external time computation tool to transform this information into the exact datetime [”2024-09-06 18:00:00”].
@@ 4.3. Progressive Order Planning
In this way, the assistant ensures reliable spatiotemporal interpretation and order planning, making it well-suited for real-world ride-hailing scenarios. Please refer to Appendix A.2.1 and A.3 for more detailed prompt template and tool information.
@@ 5.1. Multi-Type Replier Construction
In practical implementation, we provide the description of each dialog policy in the prompt to guide DiMA select the most suitable replier for response generation. To enhance response accuracy, for each dialog replier, we provide task descriptions along with demonstration examples in the prompt. Detailed prompt structures can be found in Appendix A.2.2 to Appendix A.2.5 .
@@ 8. Experiments
We conduct extensive experiments to evaluate the proposed method in both of online and offline environment. We aim to answer the following research questions: RQ1: How does DiMA perform in the online environment? RQ2: How does DiMA perform compared with existing agent frameworks? RQ3: How do different components affect the assistant’s performance? RQ4: How about the cost-effectiveness of DiMA?
@@ 8.1.4. Evaluation Protocol
In this work, we use both human and GPT evaluation.
Human Evaluation. We employ an annotation team including 7 experts, each of them has a bachelor’s degree or above in an IT-related field and possess work experience as an algorithm engineer at AI companies. Before engaging in annotation work, they received more than one week of ride-hailing annotation training. This training helps them how to learn to evaluate the correctness of trip orders and assistant responses according to serious product requirements. We provide each annotator with an average hourly wage of 7.25 dollars. Each record will be annotated by an annotator, and then another quality inspector will double-check whether the labeling is correct.
GPT Evaluation. We also propose an LLM-based evaluation method, which is widely adopted in many recent studies (Zheng et al., 2023). In this work, we use GPT-4o as the evaluator. For the trip order, we prompt GPT-4o to judge if the current order information is TRUE or FALSE. For the assistant response, we prompt GPT-4o to provide a score (from 0 to 5), and the response with a score of more than 4 will be considered as TRUE. The GPT evaluation method has also been used for annotation team to make a pre-labeling.
@@ 8.4. Ablation Study (RQ3)
Finally, we investigate the performance variation when employing continual fine-tuning. As can be seen in Figure 4, DiMA achieves consistent improvement across four metric from May 9th to May 24th, demonstrating the effectiveness of continual fine-tuning technique. The above results verify the effectiveness of important modules in DiMA. We provide data statistics over time in Appendix A.1, and an case study in Appendix A.6 for more illustrative analysis.
Figure 4. Comparison of performance over time. We fine-tune DiMA using collected data up to the latest date.
8.5. Cost-Effectiveness Analysis (RQ4)
We further discuss the cost-effectiveness of DiMA. Specifically, we replace the backbone of every module in DiMA with size-varying LLMs to analyze potential variants in accuracy and latency. As shown in Figure 5, we select, Llama3.1-405B, Mistral-123B (et al, 2023), Qwen2.5-72B, Yi-1.5-34B (Young et al., 2024), Qwen2.5-32/14/7B (Bai et al., 2023), Mistral-7B and QwQ-32B (Team, 2024) to replace the module backbones used in DiMA. It can be observed that DiMA achieves comparable performance compared with replacing backbone with much larger LLM, while the inference latency was maintained at 3.89 seconds. While Qwen2.5-72B, Mistral-123B, and Llama3.1-405B could outperform DiMA by a small margin, they spend an additional amount of latency, respectively 0.72 times (3.89s-¿6.72s), 3.17 times (3.89s-¿16.21s) and 5.47 times more (3.89s-¿25.17s). This renders them suboptimal for deployment in an online serving environment. Additionally, DiMA’s latency is higher than smaller models like Yi1.5-34B, Qwen2.5-14B, and Mistral-7B, but their performance across four metrics is significantly lower, reducing their competitiveness. Overall, the results demonstrate the advantages of DiMA’s cost-aware LLM configuration mechanism in deployment.
@@ 8.5. Cost-Effectiveness Analysis (RQ4)
9. Related Work
Task-oriented dialog systems. Task-oriented dialog (ToD) systems facilitate users in achieving specific objectives through natural language interactions, such as making reservations or inquiries (Qin et al., 2023). Numerous methodologies (Hosseini-Asl et al., 2020; He et al., 2022; Yang et al., 2021; Hudeček and Dušek, 2023; Chung et al., 2023) have been proposed to enhance ToD performance. However, existing ToD frameworks lack specialization for open-world ride-hailing scenarios. The dynamic nature of ride-hailing interactions—spanning order creation, modification, and cancellation—necessitates a tailored solution that current ToD approaches fail to provide.
LLM-based assistants. The development of domain-specific intelligent assistants powered by Large Language Models (LLMs) (Dong et al., 2023) has gained significant traction in both academia and industry. LLMs enable automated planning and execution of complex industrial tasks across various domains. For instance, Mind2Web (Deng et al., 2024), WebGPT (Gur et al., 2023), and AutoWebGLM (Lai et al., 2024) have been introduced as LLM-powered web navigation assistants. LLMPA (Guan et al., 2023) serves as a virtual assistant for travel planning within the Alipay App, while Flowris (Sun et al., 2023) and GRILLBot (Fischer et al., 2024) function as intelligent assistants for data management and multimodal conversation, respectively. Despite these advancements, no LLM-based assistant has been designed specifically to automate ride-hailing services, highlighting a gap in existing research.

### v2

@@ Abstract.
On-demand ride-hailing services like DiDi, Uber, and Lyft have transformed urban transportation, offering unmatched convenience and flexibility. In this paper, we introduce DiMA, an LLM-powered ride-hailing assistant deployed in DiDi Chuxing. Its goal is to provide seamless ride-hailing services and beyond through a natural and efficient conversational interface under dynamic and complex spatiotemporal urban contexts. To achieve this, we propose a spatiotemporal-aware order planning module that leverages external tools for precise spatiotemporal reasoning and progressive order planning. Additionally, we develop a cost-effective dialogue system that integrates multi-type dialog repliers with cost-aware LLM configurations to handle diverse conversation goals and trade-off response quality and latency. Furthermore, we introduce a continual fine-tuning scheme that utilizes real-world interactions and simulated dialogues to align the assistant’s behavior with human prefered decision-making processes. Since its deployment in the DiDi application, DiMA has demonstrated exceptional performance, achieving $93\%$ accuracy in order planning and $92\%$ in response generation during real-world interactions. Offline experiments further validate DiMA’s capabilities, showing improvements of up to $70.23\%$ in order planning and $321.27\%$ in response generation compared to three state-of-the-art agent frameworks, while reducing latency by 0.72×0.72$\times$0.72 × to 5.47×5.47$\times$5.47 ×. These results establish DiMA as an effective, efficient, and intelligent mobile assistant for ride-hailing services.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/25/08doi: 10.1145/3711896.3737208ccs: Computing methodologies Natural language generation
@@ 1. Introduction
On-demand ride-hailing services like DiDi, Uber, and Lyft have significantly transformed urban transportation systems and human mobility patterns (Xu et al., 2018; Lyu et al., 2024; Han et al., 2024). These platforms have introduced a paradigm shift in how people access transportation, offering unparalleled convenience, efficiency, and flexibility (Feng et al., 2021; Liu et al., 2022). As one of the largest ride-hailing platforms, the DiDi mobile application processes over ten billion user requests daily, encompassing tasks such as destination selection, travel time estimation, ride-hailing, re-routing, and feedback. These interactions span all stages of the ride-hailing process including pre-ride, in-ride, and post-ride (S. Alpers et al., 2020). Despite their transformative impact, existing mobile applications still rely heavily on manual inputs like typing, clicking, and scrolling, making the process time-intensive and often inefficient (Hsu and Chen, 2023).
Figure 1. The conversational interface in DiDi Chuxing App. Given a user query issued on 2024-08-28, (a) DiMA proactively guides the user to complete necessary spatiotemporal trip information, and (b) automatically schedules a future ride-hailing order from Guangdong Museum East Gate to Terminal T2 of Guangzhou Baiyun International Airport.
However, building an LLM-based ride-hailing assistant is a non-trivial task due to the following three challenges. (1) Spatiotemporal Travel Intention Understanding. The ride-hailing process relies heavily on understanding the spatiotemporal trip information (Wang et al., 2024) such as destination and departure time. LLMs often struggle with spatial reasoning (Xie et al., 2024) and temporal calculations (Su et al., 2024). For instance, when issuing a query shown in Figure 1, general LLMs may fail to derive the exact date of next Friday and neglect the user’s current location as the start location. As reported in Figure 2(a), our quantitative studies validate that even state-of-the-art LLMs like GPT-4o and Qwen2.5-72B achieve less than 80% accuracy in spatiotemporal travel intention understanding, therefore limiting their applicability for handling real-world ride-hailing requests. (2) Proactive Order Planning. Ride-hailing involves multiple factors (e.g., waiting time, willing car type, destination) for matching passengers and drivers, which may require multi-turn conversations. Moreover, during a conversation, users may cancel confirmed orders, request orders with infeasible conditions (e.g., unrealistically low costs), or engage in ride-hailing policy inquiry. How to proactively and naturally guide users through the order planning process also poses a significant challenge. (3) Cost-Effective Response Generation. The assistant should generate responses adapted to varying conversation goals and order statuses. For instance, guiding users to place an order often requires formulaic language responses, handling infeasible request needs clarification, whereas answering ride-hailing policy queries may involve retrieving DiDi’s knowledge base. As shown in Figure 2(b), in general, larger LLMs tend to be more versatile to handle complex scenarios but result in higher response latency. Generating accurate responses within strict online service latency is another critical challenge.
@@
Since May 2024, DiMA [^1] Our project is released at https://github.com/usail-hkust/DiMA. has been deployed in the DiDi application, as depicted in Figure 1. Two months’ online experiments in production demonstrate that DiMA achieves 93% accuracy in order planning and 92% in response generation. We further conduct offline experiments on synthetic datasets in Beijing and Shanghai. The results validate that DiMA significantly outperforms three state-of-the-art agent frameworks on order planning and response generation tasks, with improvements up to 70.23% and 321.27%, respectively. Moreover, the results show that DiMA maintains comparable performance to models like Qwen2.5-72B, Mistral-123B, and Llama3.1-405B, while reducing latency by 0.72x to 5.47x.
Our contributions are summarized as follows: (1) We propose the first LLM-based assistant deployed in DiDi mobile application, offering a new interaction alternative for ride-hailing; (2) We develop a spatiotemporal-aware tool-augmented order planning module and a cost-effective dialogue system to enhance accuracy and efficiency; (3) We introduce a continual fine-tuning scheme to ensure sustained proactive order planning guidance aligned with human decision processes; (4) Extensive experiments validate DiMA’s effectiveness and demonstrate its advantages against state-of-the-art LLM agent frameworks in both online and offline settings.
@@ 4.2. Temporal Reasoning
Similar to spatial mapping, we construct a four stage pipeline: (1) Temporal tool collection and annotation, (2) Temporal decomposition. (3) Tool selection, and (4) Tool calling. The whole process is similar with spatial mapping. Take the user query in Figure 1 again for example. Since the user’s departure time may not be explicitly specified, we follow product design to threat mentioned time as departure time. Therefore, DiMA will automatically generate structured function call Get_departure_time() with key parameters {”date”: ”1 week later, Fri”, ”time”: ”18:00:00”} to obtain user’s departure time. Then, based on the issued date ”2024-08-28”, DiMA calls time tool to transform it into the exact datetime [”2024-09-06 18:00:00”].
@@ 4.3. Progressive Order Planning
def Get_current_location(session_id):
POI_list = requests.post(session_id)
# [{displayname, lat, lng, id}, ...]
return POI_list
def POI_search(session_id, POI_name):
POI_list = requests.post(session_id, POI_name)
return POI_list
def POI_select(POI_list, Selected_POI_name):
return Selected_POI
def Route_planning_API(session_id, SrcLat, SrcLng,
ViaLat, ViaLng, DstLat, DstLng, SrcId, ViaId, DstId):
order_list = requests.post(session_id, ...)
return Order_list
def Time_tool(date, time, current_time):
date: "%Y-%m-%d" or "N week later, Mon" or "absent"
time: "%H:%M:%S" or "absent"
current_time: "%Y-%m-%d %H:%M:%S"
return date_time
Figure 4. API and tool lib.
@@
In this way, the assistant ensures reliable spatiotemporal interpretation and order planning, making it well-suited for real-world ride-hailing scenarios. Figure 4 shows tool information and please refer to Appendix A.2 for more detailed prompt template.
@@ 5.1. Multi-Type Replier Construction
In practical implementation, we provide the description of each dialog policy in the prompt to guide DiMA select the most suitable replier for response generation. To enhance response accuracy, for each dialog replier, we provide task descriptions along with demonstration examples in the prompt. Detailed prompt structures can be found in Appendix A.2.
@@ 8. Experiments
We conduct extensive experiments to evaluate the proposed method. We aim to answer the following research questions: RQ1: How does DiMA perform in the online environment? RQ2: How does DiMA perform compared with existing agent frameworks? RQ3: How do different components affect the assistant’s performance? RQ4: How does continual fine-tuning schema enhance DiMA? RQ5: How about the cost-effectiveness of DiMA?
@@ 8.1.4. Evaluation Protocol
In this work, we use both human evaluation and GPT self-evaluation. Human Evaluation. We employ an annotation team including 7 experts, each of them has a bachelor’s degree or above in an IT-related field and possess work experience as an algorithm engineer at AI companies. Before engaging in annotation work, they received more than one week of ride-hailing annotation training. This training helps them how to learn to evaluate the correctness of trip orders and assistant responses according to serious product requirements. We provide each annotator with an average hourly wage of 7.25 dollars. Each record will be annotated by an annotator, and then another quality inspector will double-check whether the labeling is correct. GPT Evaluation. We also propose an LLM-based evaluation method, which is widely adopted in many recent studies (Zheng et al., 2023). In this work, we use GPT-4o as the evaluator. For the trip order, we prompt GPT-4o to judge if the current order information is TRUE or FALSE. For the assistant response, we prompt GPT-4o to provide a score (from 0 to 5), and the response with a score of more than 4 will be considered as TRUE. The GPT evaluation method has also been used for annotation team to make a pre-labeling.
@@ 8.4. Ablation Study (RQ3)
In addition, we also investigate the cross-city transferability. As shown in Table 5, we adopted two experimental setups: (1) training on Beijing data and testing on Shanghai data, and (2) training on Shanghai data and testing on Beijing data. The experimental results demonstrate that the LLM-based ride-hailing assistant is not influenced by city-specific data, exhibiting strong scalability.
8.5. Continual Fine-Tuning Analysis (RQ4)
Finally, we investigate the performance variation when employing continual fine-tuning. We periodically monitoring system performance on real-world user data during the continual fine-tuning process, where new synthetic and real-user data are added to the fine-tuning dataset. Specifically, we evaluated the model on real-user interactions from May 9th to May 24th in our online system every 3 days. As shown in Figure 5, continual fine-tuning with the augmented dataset (both synthetic and real-user samples, not synthetic alone) resulted in stable and consistent performance improvements throughout this period across all four metrics, without degrading its existing performance. This confirms that the simulator-generated data effectively complements real-world samples and supports the model’s adaptation to diverse, realistic, and evolving user behavior patterns. We provide data statistics over time in Appendix A.1, and an case study in Appendix A.4 for more illustrative analysis.
8.6. Cost-Effectiveness Analysis (RQ5)
We further discuss the cost-effectiveness of DiMA. Specifically, we replace the backbone of every module in DiMA with size-varying LLMs to analyze potential variants in accuracy and latency. As shown in Figure 6, we select, Llama3.1-405B, Mistral-123B (et al, 2023), Qwen2.5-72B, Yi-1.5-34B (Young et al., 2024), Qwen2.5-32/14/7B (Bai et al., 2023), Mistral-7B and QwQ-32B (Team, 2024) to replace the module backbones used in DiMA. It can be observed that DiMA achieves comparable performance compared with replacing backbone with much larger LLM, while the inference latency was maintained at 3.89 seconds. While Qwen2.5-72B, Mistral-123B, and Llama3.1-405B could outperform DiMA by a small margin, they spend an additional amount of latency, respectively 0.72 times (3.89s-¿6.72s), 3.17 times (3.89s-¿16.21s) and 5.47 times more (3.89s-¿25.17s). This renders them suboptimal for deployment in an online serving environment. Additionally, DiMA’s latency is higher than smaller models like Yi1.5-34B, Qwen2.5-14B, and Mistral-7B, but their performance across four metrics is significantly lower, reducing their competitiveness. Notably, the recently proposed large reasoning model (e.g., QwQ-32B (Team, 2024)) failed to show significant advantages, primarily due to its poor instruction-following ability and high reasoning latency. Overall, the results demonstrate the advantages of DiMA’s cost-aware LLM configuration mechanism in deployment.
Table 5. Study of cross-city transferability of DiMA on Beijing (BJ) and Shanghai (SH).
[…]
ROA RRA SOA SRA ROA RRA SOA SRA
BJ $\to$ SH 98.03 98.92 95.96 96.81 83.95 96.13 63.20 93.13
SH $\to$ BJ 97.73 96.36 93.26 90.03 83.36 97.21 59.09 94.32
9. Related Work
This work builds upon task-oriented dialog systems and LLM-based assistant. We review representative work in both areas below.
9.1. Task-oriented Dialog Systems
Task-oriented dialog (ToD) systems facilitate users in achieving specific objectives through natural language interactions, such as making reservations or inquiries (Qin et al., 2023). In recent years, numerous methodologies (Hosseini-Asl et al., 2020; He et al., 2022; Yang et al., 2021; Hudeček and Dušek, 2023; Chung et al., 2023) have been proposed to enhance ToD system performance. However, existing ToD frameworks lack specialization for open-world ride-hailing scenarios. The dynamic nature of ride-hailing interactions—spanning order creation, modification, and cancellation—necessitates a tailored solution that current ToD approaches fail to provide.
Figure 5. Comparison of DiMA’s performance over the time. We periodically (every three days) fine-tune DiMA using collected data up to the latest date.
@@ 9.1. Task-oriented Dialog Systems
9.2. LLM-based Assistants
The development of domain-specific intelligent assistants powered by Large Language Models (LLMs) (Dong et al., 2023) has gained significant traction in both academia and industry. LLMs enable automated planning and execution of complex industrial tasks across various domains. For instance, Mind2Web (Deng et al., 2024), WebGPT (Gur et al., 2023), and AutoWebGLM (Lai et al., 2024) have been introduced as LLM-powered web navigation assistants. LLMPA (Guan et al., 2023) serves as a virtual assistant for travel planning within the Alipay App, while Flowris (Sun et al., 2023) and GRILLBot (Fischer et al., 2024) function as intelligent assistants for data management and multimodal conversation, respectively. Despite these advancements, no LLM-based assistant has been designed specifically to automate ride-hailing services, highlighting a gap in existing research.